OpenStudy (anonymous):

F(x)= x^2+8x Express the quadratic function in standard form

zepdrix (zepdrix):

Hmm so we have a parabola, I think standard form looks something like this if I remember correctly: \[f(x)=(x-h)^2 + k\] So we need to complete the square, then we'll have a left over quantity (k). Make sense? :D

OpenStudy (anonymous):

whats my x and whats my h?

zepdrix (zepdrix):

It'll make a little more sense once you see the answer :) First step is to complete the square, So we take half of the b term (half of 8) and square it.

OpenStudy (anonymous):

x^2+8x+16

zepdrix (zepdrix):

\[f(x)=x^2+8x+16-16\] Ok good! :) but we can't just add 16 willy nilly! we have to balance the equation like this.

OpenStudy (anonymous):

thats the same as x^2+8x

zepdrix (zepdrix):

Yes, but we wont let the 16's cancel out, we're going to do something with them ^^

zepdrix (zepdrix):

x^2+8x+16 is a perfect square, what does it simply into?

OpenStudy (anonymous):

(x+4)^2

zepdrix (zepdrix):

\[f(x)=(x^2+8x+16)-16\]\[f(x)=(x+4)^2-16\] Ok very good :) and we still have the 16 left over right? Is it a litte easier to recognize your H and K now?

zepdrix (zepdrix):

Uh ohhhh, Brent confused? :D

OpenStudy (anonymous):

h=-4 k=-16

zepdrix (zepdrix):

Ok good good :D And what is significant about that point is the fact that it tells us where our vertex is located! That center point from which the parabola grows in 2 directions :D

OpenStudy (anonymous):

i need to put it in standard form and find the vertex and x/y intercepts

zepdrix (zepdrix):

\[f(x)=(x+4)^2-16\] Ok so we've gotten it into standard form. You identified h and k. (h,k) is your vertex. Now to find intercepts, do the following: For y-intercepts, plug in x=0 and solve for f(x). For x-intercepts, plug in f(x)=o and solve for x (you should get 2 values for this one).

OpenStudy (anonymous):

so then there is a y-int at 0

zepdrix (zepdrix):

the y-axis is intercepted (crossed by our function) when x=0. When we plug x=0 into our function, we get: \[f(x)=(0+4)^2-16\]\[f(x)=0\] Ah so we do! :) Yes good job.

OpenStudy (anonymous):

what about the x?

zepdrix (zepdrix):

We'll plug 0 in for the other variable and solve for x.

OpenStudy (anonymous):

X^2+2*x*4+4^2-4^2, so (x+4)^2-16

zepdrix (zepdrix):

\[0=(x+4)^2-16\] Solve for x from here. Important note, when you take the square root of both sides, you'll actually get 2 values for (x+4)! :D Ok i gotta go, test tomorrow :D gotta get some sleep. Good luck with the rest of the problem Brent!

OpenStudy (anonymous):

its 0!

OpenStudy (anonymous):

so x has an intercept at zero too